OpenStudy (anonymous):

Hey, I need help with the method of solving the series 2^n/(3^(n-1)) from n=1 to infinity

OpenStudy (anonymous):

Okay, so you have: \[\sum_{n=1}^{\infty}\frac{2^n}{3^{n-1}}\]?

OpenStudy (anonymous):

yup, ps nice pic haha

OpenStudy (anonymous):

easy

OpenStudy (anonymous):

= 3 times sum of (2/3)^n

OpenStudy (anonymous):

then S= a/(1-r)

OpenStudy (anonymous):

everyone wants a piece of this one

OpenStudy (anonymous):

= (3 x (2/3) ) / (1/3)

OpenStudy (anonymous):

= 6

OpenStudy (anonymous):

Haha, thank you. So what you need to do is get in geometric series form. You know that geometric series always start with a 1 but if you plug in n=1, you notice you get 2? So factor out a 2 to rewrite it. Giving: \[\sum_{n=1}^{\infty}2(\frac{2}{3})^{n-1}\]. So when you plug in n=1 you get a 1. So the series converges to: \[\frac{2}{1-\frac{2}{3}}\]. Since |2/3|<1. Which is 6.

OpenStudy (anonymous):

Well a 2(1) :P Since the two can be pulled out of the entire summation.

OpenStudy (anonymous):

thanks, didn't see it as a geometric series, thanks everyone